Five different digits 7 3 6 2 5 are given. Out of these digits, 4 digits numbers are formed. Find the no. of 4-digit numbers formed which have at least one digit repeated.
Answer by Edwin McCravy(20054)

First find the number of four-digit numbers.

5 ways to pick the 1st digit.
5 ways to pick the 2nd digit.
5 ways to pick the 3rd digit.
5 ways to pick the 4th digit.


That's 5×5×5×5 = 54 = 625 four digit numbers altogether.

We need to subtract from that the number of four-digit numbers
which have no digit repeated.

5 wasy to pick the 1st digit.
4 ways remaining to pick the 2nd digit.
3 ways remaining to pick the 3rd digit.
2 ways rewmaining to pick the 4th digit.

That's 5×4×3×2 = 5! = 120 four digit numbers with no digits repeated.

Answer: 625-120 = 505.

Edwin